1 mole BaF2 +2 mole H2SO4 gives rise to resulting mixture it will be neutralized by -
a) 1 mol of KOH
b)  2 mol of H2SO4
c)  4 mol KOH
d)  2 mol of KOH

Grade:12

4 Answers

Arun
5 years ago
The base OH- exactly neutralises mole of H+ion, for example 1 mol of OH- will neutralise 1 mol of H+ ion.

So as per questionnaire 2 moles of H+ ion from H2SO4 will reacts with 2 moles of OH- for neutralisation.

Hence answer is Option D) 2 moles KOH.
 
I am not sure about my answer
Ayush Mittal
5 years ago
It is a multiple correct question.The answer given is option B and option C. Please think about it again. It would be much grateful.
Regards
Thagaval medai
5 years ago
BaF2 +H2SO4=BaSO4+2HF
(BaSO4+2HF)+X=products (neutralization)
Mole*n factor=equivalence
BaSO4:-
1*2=2
2HF:-
2*1=2
Equivalence(BaSO4+2HF)=Equivalence (X)(neutralization rxn)
2+2=moles*n factor of X
1)KOH
4=n*1
n=4(no.of moles)
Hence 4KOH
2)Ca(OH)2
4=n*2
n=2(no.of moles)
Hence 2Ca(OH)2
Thus 4 KOH & 2Ca(OH)2
B
5 years ago
BaF2+ H2SO4= BaSO4+ 2HF  
          
1 mol each of the above reactants are consumed 
and 2 moles of HF is formed.
Now the mixture has 1 mol of H2SO and 2 moles
of HF.
For neutralisation 4 moles of OH- will be required.
Hence, B (which should be 2 moles of Ca(OH)
and C will be the right answer.